# MAPO: Momentum-Aided Gradient Descent Prompt Optimization

### v2

@@ Abstract
Momentum-Aided Prompt Optimization (MAPO) enhances the efficiency and efficacy of prompt optimization for Large Language Models (LLMs). Building on ProTeGi (Pryzant et al., 2023), MAPO uses positive natural language "gradients" and a momentum-based extension to refine prompts effectively. By tracking gradient history, MAPO avoids local minima and oscillations. It also utilizes beam search and an Upper Confidence Bound (UCB) algorithm for balanced candidate expansion and selection. Benchmark testing shows that MAPO achieves faster convergence time with fewer API calls and higher F1 scores than ProTeGi, proving it as a robust and scalable solution for automated prompt engineering in LLMs.
MAPO: Momentum-Aided Gradient Descent Prompt Optimization
Anthony Cui           Pranav Nandyalam           Ethan Cheung           Kevin Zhu Algoverse AI Research acui@overlake.org, pranavrajnandyalam@gmail.com, kevin@algoverse.us
@@ 1 Introduction
Large Language Models (LLMs) have gained significant attention since the release of ChatGPT OpenAI (2022), leading to the development of new prompting techniques that have greatly improved LLM performance (Schulhoff et al., 2024). However, prompts can still be unclear, biased, or incomplete, limiting LLM capabilities (Sahoo et al., 2024). Prompt engineering has become critical, but current methods often require manual adjustments, making them time-consuming, error-prone, and constrained by human limitations (Lin et al., 2024). This highlights an increasing need for an automated system to improve prompt quality.
Recent work has explored implementing traditional machine learning algorithms into a natural language format, with one of the first being ProTeGi’s “Automatic Prompt Optimization with ‘Gradient Descent’ and Beam Search” (Pryzant et al., 2023). While ProTeGi introduced an innovative framework, it has limitations such as excessive API calls, resource consumption, and underutilizing the strengths of prompts.
@@
We introduce Momentum-Aided Prompt Optimization (MAPO), a method that extends ProTeGi by using positive natural language “gradients” with momentum to automate prompt refinement. Gradients are generated from correct examples in a minibatch, guiding the LLM to refine prompts in a consistent semantic direction. Beam search expands the candidate pool, and the best-arm identification algorithm, using UCB bandits, selects the top prompts for further evaluation.
MAPO improves on ProTeGi by speeding up convergence and reducing resource use through momentum-based adjustments. By tracking gradient history, MAPO avoids local minima and oscillations. In a case study, MAPO demonstrates faster runtime, higher F1 scores, and fewer API calls, providing a scalable solution for automated prompt engineering in LLMs.
@@ 2 Related Works
Prompt Engineering. In this work, we draw upon existing prompt engineering techniques and focus on incorporating optimization algorithms into our framework to enhance the effectiveness of prompt optimization. There is an increasingly diverse set of general frameworks that previous works have focused on: LLM optimization ((Pryzant et al., 2023); (Zelikman et al., 2023); (Fernando et al., 2023); (Zhou et al., 2022); (Yang et al., 2023)), reinforcement learning ((Ma et al., 2023); (Zhang et al., 2022); (Deng et al., 2022)), and in-context learning (Shum et al., 2023). However, these approaches are generally not feasible when there is no architectural information introduced and only an API is provided to the LLM. More specifically, we base most of our work on improving automatic prompt engineering techniques with LLM gradient-based methods ((Shin et al., 2020) (Pryzant et al., 2023)). Though, many of the current methods involving some sort of iterative refinement technique, such as APE (Zhou et al., 2022) and ProTeGi (Pryzant et al., 2023) all face similar struggles with cumulative costs of running their programs.
@@ 3.1 Momentum-Aided Prompt Optimization
First, the current prompt $\mathbf{p}$ is evaluated based on a minibatch of training data to obtain randomly sampled strings $\mathbf{s}$, which contain correct LLM predictions generated relative to the correct labels. We then provide the LLM with a static prompt $\tau$ to generate numerous positive “gradients” $\nabla\mathbf{p}$ in natural language, praising the current prompt $\mathbf{p}$ using the sampled strings $\mathbf{s}$. Our “gradients” $\nabla\mathbf{p}$ are the natural language outputs of the LLM’s continuation of static prompt $\tau$. In traditional machine learning, gradient descent uses numerical gradients, representing a vector in parameter space where a model can improve or worsen; by contrast, our textual gradients $\nabla\mathbf{p}$ represent directions in semantic space (Pryzant et al., 2023). We use another static prompt $\alpha$ to apply these gradients to the initial prompt $\mathbf{p}$, allowing us to move along the same semantic direction as the positive textual gradients, refining and improving the initial prompt.
@@ 3.2 Expansion and Selection
Our method employs beam search to explore the space of prompt variations generated during optimization. In each round, new candidate prompts $\mathbf{p}_{k}$ are created from the top $k$ best performing prompts from the previous round by iterating through the prompt optimization process described in Section 3.1. After each round, less promising candidates are pruned, and only the top $k$ prompts are retained for further gradient-based improvements, evaluated by a scoring function that assesses how well they meet our predefined objectives, such as F1 score.
@@
We also record the gradients from static prompt $\alpha$, used to generate the top $k$ candidates in each round, to incorporate our novel momentum extension into the natural language gradient descent. Drawing on the physics intuition of momentum, traditional gradient descent uses this extension to improve stability and convergence, helping the model avoid oscillations and escape local minima, thereby reaching global minima more efficiently. Analogously, our method maintains a history of past gradients, guiding the movement of the initial prompt $\mathbf{p}$ in each beam search round through semantic space, helping it converge on the optimal prompt rather than just an incremental improvement. A single positive gradient is randomly sampled from a pool of all the gradients used to generate the top $k$ prompts in each beam search round, representing the positive gradient history. This gradient history is then incorporated into our static prompts $\tau$ and $\alpha$ as textual momentum, guiding the LLM to generate new gradients $\nabla\mathbf{p}$ and new prompt candidates $\mathbf{p}_{c}$, allowing the initial prompt $\mathbf{p}$ to “roll down the hill” faster, i.e., achieve a much faster rate of convergence during prompt optimization.
In our implementation, we utilize the same Upper Confidence Bound (UCB) Bandits Selection algorithm as ProTeGi (Pryzant et al., 2023), which is employed each beam search expansion to evaluate the candidate prompts. UCB has proven to be the strongest best arm identification algorithm for maximizing test metrics such as F1 score, outperforming other algorithms like UCB-E, Successive Rejects, and Successive Halving (Pryzant et al., 2023). UCB effectively balances exploration—searching for better options—and exploitation—selecting the best-performing candidates.
Figure 2: Test performance (F1) versus Time for Liar and Ethos datasets.
Figure 3: Test performance (F1) versus Total API Calls for Liar and Ethos datasets.
Figure 4: Test performance (F1) versus Beam Search Round for Liar and Ethos datasets.
@@ 4.1 Setup
Our experimental setup closely follows that of ProTeGi, allowing for a direct comparison between our extension method and their baseline. We use 200 randomly sampled data points as the test set, retaining most hyperparameters from ProTeGi’s configuration, including a temperature of 0, a minibatch size of 64, and 6 rounds of beam search with a selection size of 4. In each iteration (as outlined in Section 2.1), we generate 2 positive gradients using 3 randomly sampled correct examples from the minibatch, resulting in fewer gradients than ProTeGi’s setup, which uses 4 negative gradients. This trade-off improves runtime efficiency while handling the added complexity of gradient history. The primary evaluation metric is the F1 score, and results reflect the highest score among the top k beam search candidates, averaged over three trial runs to account for variability. Unless otherwise stated, all experiments use the October 2024 release of GPT-3.5-turbo.
@@ 4.1.1 Baseline
ProTeGi. Developed by Pryzant et al. (2023), ProTeGi employs natural language gradient descent with negative gradients from incorrect example sampling to refine prompts. It iteratively applies these gradients to address prompt weaknesses, expanding the candidate pool using Monte-Carlo sampling to generate paraphrased versions with synonyms or semantically similar variations. This ensures candidate diversity while guiding the optimization process.
@@ 4.1.2 Benchmarks
ProTeGi has been evaluated on benchmark datasets, such as the Liar Dataset Wang (2017) for fake news detection and the Ethos Dataset Mollas et al. (2022) for hate speech detection, which test its ability to handle diverse tasks. Our method will be evaluated on the same benchmarks for comparison. Despite utilizing the publicly available code from ProTeGi to replicate the baseline, we were unable to reproduce the level of metric performance for their method on the Liar Dataset as reported in the original paper.
@@ 4.2 Analysis
The initial prompt itself has baseline F1 scores of 0.475 and 0.9 for Liar and Ethos, respectively, evaluated on our test set, serving as the starting point for optimization.
Efficiency. Figures 2, 3, and 4 show the significant efficiency gains of MAPO over ProTeGi. In Figure 2 for Liar, ProTeGi takes approximately 735 seconds to reach an F1 score of 0.58, while MAPO achieves the same score in only 285 seconds. For Ethos, ProTeGi takes 686 seconds while MAPO takes 136 seconds to reach the same level of performance. Overall, there is a 72.74% reduction in runtime. This is notable since ProTeGi’s runtimes can extend into hours (Pryzant et al., 2023), highlighting the resource-intensive nature of automatic prompt optimization.
Figure 3 illustrates that ProTeGi uses about 417 API calls to reach its final F1 score, whereas MAPO requires an average of 77 to reach the same performance, resulting in an average 81.53% reduction. MAPO also completes all six rounds of beam search with 105 fewer API calls on average, addressing critical computational constraints in large-scale optimization.
@@
While MAPO incurs longer processing times per iteration due to more complex prompt structures and the inclusion of positive gradient history, the overall reduction in steps and runtime demonstrates that these trade-offs do not detract from its overall efficiency.
Efficacy. Figure 4 shows that MAPO consistently outperforms ProTeGi at every beam search round for both Liar and Ethos. MAPO’s F1 score steadily increases while ProTeGi quickly converges and then plateaus or slightly declines. This results in a notable 5.37% increase in overall performance for MAPO, underscoring its superior efficacy in optimizing model outputs.
MAPO’s consistent improvement highlights its effectiveness in leveraging positive gradients and the momentum-based extension to thoroughly explore and refine the prompt search space. By incorporating momentum, MAPO maintains a consistent direction in the optimization process, helping it avoid local minima and erratic updates that hinder progress. This leads to a more robust optimization mechanism compared to ProTeGi. In contrast, ProTeGi’s performance dips suggest it struggles with optimization challenges like local minima and oscillations due to the lack of a stabilizing mechanism. MAPO’s use of momentum and positive gradient history allows it to maintain steady progress, effectively "remembering" beneficial adjustments and reducing the likelihood of stagnation. This results in more reliable convergence toward higher performance levels, demonstrating the superior efficacy of MAPO’s optimization strategy.
@@
However, there is a significant difference in convergence speed: on average, we observe a 54% decrease in convergence time when incorporating momentum into MAPO. This substantial improvement in convergence speed demonstrates that momentum effectively enhances the optimization process by smoothing the loss landscape and helping the model avoid getting trapped in local minima. This is further evidenced by our smoother test set curves depicted in Figures 2, 3, and 4, which contrast sharply with the oscillations observed in ProTeGi’s data. The lack of significant fluctuations in our graphs confirms that our method not only maximizes convergence speed but also promotes stability during training. Collectively, these findings validate the efficiency of incorporating momentum in prompt optimization, proving that our approach accelerates convergence while maintaining, or even slightly improving, the model’s peak performance.

### v3

# Momentum-Aided Natural Language Gradient Descent for Prompt Optimization

@@ Abstract
Prompt optimization is crucial for improving the output quality of Large Language Models (LLMs), but many existing methods are inefficient, requiring extensive computation and manual tuning. We propose Momentum-Aided Prompt Optimization (MAPO), which builds on ProTeGi (Pryzant et al., 2023) by incorporating positive natural language "gradients" and a momentum-based memory mechanism to refine prompts while avoiding local minima and oscillations. It also employs beam search and an Upper Confidence Bound (UCB) algorithm for balanced candidate expansion and selection. MAPO achieves faster convergence time with fewer API calls and higher performance than ProTeGi, demonstrating its effectiveness as a robust and scalable solution for automated prompt optimization in LLMs. Our code is available online at https://github.com/AnthonyCui7/momentum-aided-prompt-optimization.
Momentum-Aided Natural Language Gradient Descent for Prompt Optimization
Anthony Cuithanks: Lead Author  Pranav Nandyalam  Andrew Rufail  Ethan Cheung Aiden LeiKevin ZhuSean O’Brien Algoverse AI Research anthonycui@u.northwestern.edu, kevin@algoverse.us
@@ 1 Introduction
Large Language Models (LLMs) have gained significant attention since the release of ChatGPT OpenAI (2022), leading to the development of new prompting techniques that have greatly improved LLM performance (Schulhoff et al., 2024). While it has been shown that the prompts given to the LLM greatly affect performance (Pawlik, 2025), prompts can still be unclear, biased, or incomplete, limiting LLM capabilities (Sahoo et al., 2024). For these reasons, prompt engineering has become a critical aspect of leveraging an LLM’s capabilities. Often, current prompt engineering methods require manual adjustments by the user, making them time-consuming, error-prone, and constrained by human limitations (Lin et al., 2024). This highlights an increasing need for an automated system to improve the quality of prompts without the need for human intervention.
@@
Recent work has explored implementing traditional machine learning algorithms in a natural language format, with one of the first being ProTeGi’s “Automatic Prompt Optimization with ‘Gradient Descent’ and Beam Search” (Pryzant et al., 2023). While ProTeGi introduced an innovative framework, it has limitations, including high computational costs and resource consumption due to excessive API calls making large-scale prompt optimization impractical. Furthermore, ProTeGi does not track previous refinements, leading to oscillatory behavior and slow convergence. Ultimately, the strengths of prompts are underutilized.
We introduce Momentum-Aided Prompt Optimization (MAPO), a method that extends ProTeGi by using positive natural language “gradients” with momentum to automate prompt refinement. Gradients are suggested feedback for the current prompt generated using correct examples from a minibatch of training data, guiding the LLM to refine prompts in a consistent semantic direction. Beam search expands the generated candidate pool of improved child prompts, and the best-arm identification algorithm using UCB bandits, selects the top prompts for further evaluation. MAPO significantly improves upon ProTeGi by enhancing efficiency and effectiveness through momentum-based adjustments. Unlike ProTeGi, which treats each optimization step as an independent improvement, MAPO tracks the history of gradient updates, preventing oscillations and reducing the risk of suboptimal prompts (local minima in the semantic space). In our experimental results in Section 4, MAPO achieves a 77.9% reduction in convergence time compared to ProTeGi, 88.0% reduction in convergence API calls, as well as a 5.28% increase in peak performance, providing a scalable solution for automated prompt engineering in LLMs.
@@ 2 Related Works
Prompt Optimization. In this work, we draw upon existing prompt engineering techniques and focus on incorporating optimization algorithms into our framework to enhance the effectiveness of prompt optimization. There is an increasingly diverse set of general frameworks that previous works have focused on: LLM optimization (Pryzant et al., 2023; Zelikman et al., 2023; Fernando et al., 2023; Zhou et al., 2022; Yang et al., 2023), reinforcement learning (Ma et al., 2023; Zhang et al., 2022; Deng et al., 2022), and in-context learning (Shum et al., 2023). However, these approaches are often infeasible when there is no architectural information introduced and only an API is provided to the LLM. More specifically, we base most of our work on improving automatic prompt engineering techniques with LLM gradient-based methods (Shin et al., 2020; Pryzant et al., 2023). Though, many of the current methods involving some form of iterative refinement technique, such as APE (Zhou et al., 2022) and ProTeGi all face similar struggles with cumulative costs of running their programs.
@@ 3.1 Momentum-Aided Prompt Optimization
First, the current prompt $\mathbf{p}$ is evaluated based on a minibatch of training data to obtain randomly sampled strings $\mathbf{s}$ representing correct LLM predictions aligned with ground-truth labels. We then provide the LLM with a static prompt $\tau$ (see Appendix A) to generate numerous positive “gradients” $\nabla\mathbf{p}$ in natural language, appraising the current prompt $\mathbf{p}$ using the sampled strings $\mathbf{s}$. Our “gradients” $\nabla\mathbf{p}$ are the natural language outputs of the LLM’s continuation of static prompt $\tau$. In traditional machine learning, gradient descent uses numerical gradients, representing a vector in parameter space that points in the direction of steepest ascent of the loss function, which the algorithm counters by moving in the opposite direction to minimize the loss function and improve model performance; in contrast, our natural language gradients $\nabla\mathbf{p}$ represent directions in semantic space (Pryzant et al., 2023). We use another static prompt $\alpha$ (Appendix A) to apply these gradients to the initial prompt $\mathbf{p}$, allowing us to move along the same semantic direction as the positive natural language gradients, refining and improving the initial prompt.
@@ 3.2 Expansion and Selection
Algorithm 1 MAPO Beam Search
1: $p_{0}$: initial prompt, $b$: beam width, $r$: search depth, $m$: metric function, $G$: gradient history
2: $B_{0}\leftarrow\{p_{0}\}$
3: $G_{0}\leftarrow\emptyset$
4: for $i\leftarrow 0$ to $r-1$ do
5:      $C\leftarrow\emptyset$
6:      for all $p\in B_{i}$ do
7:          $C\leftarrow C\cup\textit{Expand}(p,G_{i})$triangleright Section 3.2
8:      end for
9:      $B_{i+1}\leftarrow\textit{Select}_{b}(C,m)$triangleright UCB Bandits
10:      $G_{i+1}\leftarrow\textit{SampleGradient}(B_{i+1})$
11: end for
12: $\hat{p}\leftarrow\textit{argmax}_{p\in B_{r}}~{}m(p)$
13: return $\hat{p}$
@@
As outlined in Algorithm 1, our method employs beam search to explore the space of prompt variations generated during optimization. In each optimization round, the LLM utilizes static prompt $\alpha$ containing sampled strings $\mathbf{s}$ and gradients $\nabla\mathbf{p}$ to edit the top $k$ best-performing prompts from the previous round, producing $c$ new candidate prompts, denoted as $\mathbf{p}_{c}$. After each round, less promising candidates are pruned, and only the top $k$ prompts are retained for further gradient-based improvements, evaluated by a scoring function that assesses how well they meet our predefined objectives, such as the F1 score (Manning et al., 2008).
In our implementation, we utilize an Upper Confidence Bound (UCB) Bandit Selection algorithm, which is applied during each beam search expansion to evaluate candidate prompts. The UCB Bandit Selection algorithm we employ is outlined in Algorithm 2 (Pryzant et al., 2023).
Algorithm 2 UCB Bandits
1: $n$ prompts $p_{1},\dots,p_{n}$, dataset $\mathcal{D}_{\mathrm{tr}}$, $T$ time steps, metric function $m$, exploration $c_{v}$
2: Initialize $N(p_{i})\leftarrow 0,\;Q(p_{i})\leftarrow 0\quad\forall i\in[1..n]$
3: for $t=1,\dots,T$ do
4:      Sample $\mathcal{D}_{\text{sample}}\subset\mathcal{D}_{\text{tr}}$ uniformly
5:      $p_{i}\leftarrow\arg\max_{p}\left\{Q_{t}(p)+c_{v}\sqrt{\frac{\log t}{N_{t}(p)}}\right\}$
6:      Observe reward $r_{i,t}=m(p_{i},\mathcal{D}_{\text{sample}})$
7:      $N_{t}(p_{i})\leftarrow N_{t}(p_{i})+|\mathcal{D}_{\text{sample}}|$
8:      $Q_{t}(p_{i})\leftarrow Q_{t}(p_{i})+\frac{r_{i,t}}{N_{t}(p_{i})}$
9: end for
10: return $\textsc{SelectTop}_{b}(Q_{T})$
@@
We also record the gradients from static prompt $\alpha$ used to generate the top $k$ candidates in each round to incorporate our novel momentum extension into natural language gradient descent. Drawing on the physics intuition of momentum, traditional gradient descent uses this extension to improve stability and convergence, helping the model avoid oscillations and escape local minima, thereby reaching global minima more efficiently (Malingan, 2023). Analogously, our method maintains a history of past gradients, guiding the movement of the initial prompt $\mathbf{p}$ in each beam search round through semantic space, helping it converge on the optimal prompt rather than just an incremental improvement. A single positive gradient is randomly sampled from a pool of all the gradients used to generate the top $k$ prompts in each beam search round, representing the positive gradient history. This gradient history is then incorporated into our static prompts $\tau$ and $\alpha$ as textual momentum, guiding the LLM to generate new gradients $\nabla\mathbf{p}$ and new prompt candidates $\mathbf{p}_{c}$, allowing the initial prompt $\mathbf{p}$ to “roll down the hill” faster, i.e., achieve a much faster rate of convergence during prompt optimization.
@@ 4.1 Setup
Our experimental setup closely follows that of ProTeGi, allowing for a direct comparison between our extension method and their baseline. We use 200 randomly sampled data points as the test set, retaining most hyperparameters from ProTeGi’s configuration, including a temperature of 0, a minibatch size of 64, and 6 rounds of beam search with a beam width of 4. Each parent prompt expands into 8 new candidate prompts, with 2 positive gradients generated from 3 randomly sampled correct examples from the minibatch, which creates fewer gradients than ProTeGi’s method of using 4 negative gradients. This trade-off improves runtime efficiency while handling the added complexity of gradient history. The primary evaluation metric is the F1 score, and results reflect the highest score among the top $k$ beam search candidates, averaged over three trial runs. All experiments use the October 2024 release of GPT-3.5-turbo unless otherwise stated.
Figure 2: Test performance (F1 Score) versus Time for Liar, Ethos, and Word-in-Context dataset.
Figure 3: Test performance (F1 Score) versus Cumulative API Calls for Liar, Ethos, and Word-in-Context dataset.
Figure 4: Test performance (F1 Score) versus Beam Search Round for Liar, Ethos, and Word-in-Context datasets.
Dataset Method/Metric GPT-3.5-Turbo GPT-4o Mini
Ethos Initial F1 0.9 0.975
MAPO 0.98 / 871 / 364 / 146 / 56 / 2 0.985 / 1325 / 364 / 96 / 19 / 1
ProTeGi 0.965 / 687 / 472 / - / - / - 0.985 / 1328 / 423 / - / - / -
Liar Initial F1 0.475 0.52
MAPO 0.64 / 904 / 364 / 260 / 98 / 3 0.625 / 1414 / 364 / - / - / -
ProTeGi 0.58 / 737 / 472 / - / - / - 0.64 / 1362 / 399 / - / - / -
WiC Initial F1 0.52 0.665
MAPO 0.66 / 937 / 364 / 73 / 16 / 1 0.73 / 1252 / 364 / 533 / 157 / 3
ProTeGi 0.635 / 755 / 472 / - / - / - 0.73 / 1119 / 406 / - / - / -
Table 1: Comparison of MAPO (ProTeGi) and baselines across tasks and models. Each cell shows: Final F1 Score / Total Time (s) / Total API Calls / Convergence Time (s) / Convergence API Calls / Convergence Optimization Steps. The row Initial F1 reports performance of the initial prompt before optimization.
Liar Dataset
F1 Score Conv. Time (s) Total Time (s)
Positive Only (MAPO) 0.66 285 903
Negative Only 0.60 303 1228
Both Positive & Negative 0.65 275 1467
Table 2: Comparison of Past Gradient Methods on the Liar Dataset. Convergence time represents time to reach peak ProTeGi performance for Liar
@@ 4.1.1 Baseline
ProTeGi. Developed by Pryzant et al. (2023), ProTeGi employs natural language gradient descent with negative gradients from incorrect example sampling to refine prompts. It iteratively applies these gradients to address prompt weaknesses, expanding the candidate pool using Monte Carlo sampling to generate paraphrased versions with synonyms or semantically similar variations. This ensures candidate diversity while guiding the optimization process.
@@ 4.1.2 Benchmarks
ProTeGi has been evaluated on benchmark datasets, such as the Liar Dataset (Wang, 2017), which focuses on detecting fake news by classifying claims as true or false, and the Ethos Dataset (Mollas et al., 2022), which is designed for identifying hate speech in online conversations. Our method will be evaluated on these same benchmarks for comparison. Additionally, we evaluate on the Word-in-Context dataset (Pilehvar and Camacho-Collados, 2018), which tests models on their ability to understand and represent the meanings of certain words based on the context in which they appear, to further expand the range of NLP tasks tested. Despite utilizing the publicly available code from ProTeGi to replicate the baseline, we were unable to reproduce the level of metric performance for their method on the Liar Dataset as reported in the original paper.
@@ 4.2 Analysis
Efficiency. Figures 2, 3, and 4 show the significant efficiency gains of MAPO over ProTeGi in terms of convergence time for MAPO to reach the same peak F1 score for ProTeGi. Overall, there is an average 77.9% reduction in convergence time. This is notable since ProTeGi’s runtimes can extend into hours (Pryzant et al., 2023), highlighting the resource-intensive nature of automatic prompt optimization.
Figure 3 illustrates that ProTeGi uses about 417 API calls to reach its final F1 score, whereas MAPO requires an average of 56 to reach the same performance, resulting in an average 88.0% reduction. MAPO also completes all six rounds of beam search with 105 fewer API calls on average, addressing critical computational constraints in large-scale optimization.
@@
We attribute MAPO’s efficiency primarily to its use of momentum, which allows it to incorporate feedback from previous optimization steps when refining prompts. Instead of treating each update independently, MAPO builds on prior successful changes, resulting in more consistent progress, fewer wasted iterations, and significantly faster convergence across multiple metrics. The use of positive gradients further constrains the search space to prompt edits supported by observed success, rather than encouraging arbitrary variation. Together, these design choices enable a more focused and informed optimization process, reducing the number of steps and API calls required to reach strong performance.
While MAPO does incur longer processing times per iteration due to more complex prompt structures and the inclusion of positive gradient history, the overall reduction in steps and runtime demonstrates that these trade-offs do not detract from its overall efficiency.
Liar Dataset
Time (s) F1 Score
No Momentum 390 0.64
With Momentum 285 0.64
Ethos Dataset
Time (s) F1 Score
No Momentum 407 0.98
With Momentum 77 0.985
Table 3: Comparison of Momentum Ablation on Liar and Ethos Datasets: Peak F1 Score and Convergence Time to Reach Peak ProTeGi Performance
Math Dataset Accuracy
GSM8K AGIEval-Math
MAPO 95.5 75.3
Original prompt 94.5 74.1
Table 4: Comparison of MAPO’s performance on math datasets.
Efficacy. Figure 4 shows that MAPO consistently outperforms ProTeGi at every beam search round for both Liar and Ethos. MAPO’s F1 score steadily increases while ProTeGi quickly converges and then plateaus or slightly declines. This results in a notable 5.28% average increase in peak performance for MAPO.
MAPO’s consistent improvement highlights its effectiveness in leveraging positive gradients and the momentum-based extension to thoroughly explore and refine the prompt search space. By incorporating momentum, MAPO helps it avoid local minima and erratic updates that hinder progress. This leads to a more robust optimization mechanism compared to ProTeGi. In contrast, ProTeGi’s performance dips suggest it struggles with optimization challenges like local minima and oscillations due to the lack of a stabilizing mechanism. MAPO’s use of momentum and positive gradient history allows it to maintain steady progress, effectively "remembering" beneficial adjustments and reducing the likelihood of stagnation. This results in more reliable convergence toward higher performance levels, demonstrating the superior efficacy of MAPO’s optimization strategy.
Evaluation Across Models. Focusing on the performance of GPT-4o mini in Table 1, the model consistently demonstrates a strong initial F1 score, likely attributable to its advanced architecture. While ProTeGi yields competitive F1 scores with GPT-4o mini, MAPO achieves notably faster convergence in terms of optimization steps (1 or 3 steps) and convergence API calls (e.g., Ethos: 19 API calls for MAPO to reach ProTeGi performance with 423 API calls). This consistency of improved efficiency reinforces MAPO’s utility as a scalable and adaptable prompt optimization framework beyond the original GPT-3.5-Turbo setting.
Gradient Ablation. As shown in Table 2, the method using only positive gradients achieves the highest F1 score with the shortest total time, demonstrating its efficiency and effectiveness. The method combining both positive and negative gradients has the fastest convergence time, but its F1 score is slightly lower, and it requires the longest total time. This suggests that while the combined method offers faster convergence, it sacrifices some performance. On the other hand, the method using only negative gradients performs the worst across all metrics. These results highlight the importance of positive gradients for optimizing both accuracy and efficiency, while the combined method excels in rapid convergence scenarios. Although it takes longer to converge, the positive-only method significantly reduces overall runtime, as well as a slightly higher peak F1 score. Overall, we conclude that the benefits of a significantly shorter total time and a higher F1 score outweigh the slightly higher convergence time.
@@
However, there is a significant difference in convergence speed: on average, we observe a 54% decrease in convergence time when incorporating momentum into MAPO. This substantial improvement in convergence speed demonstrates that momentum effectively enhances the optimization process by smoothing the loss landscape and helping the model avoid getting trapped in local minima, an issue that ProTeGi struggles with Pryzant et al. (2023). This is further evidenced by our smoother test set curves depicted in Figure 2, 3, and 4, which contrast sharply with the oscillations observed in ProTeGi’s data. The lack of significant fluctuations in our graphs confirms that our method not only maximizes convergence speed but also promotes stability during training. Collectively, these findings validate the efficiency of incorporating momentum in prompt optimization, proving that our approach accelerates convergence while maintaining, or even slightly improving, the model’s peak performance.
Mathematical Reasoning. Since MAPO is a general prompt optimization algorithm, it can be easily generalized to many other tasks, not just binary classification. We test MAPO on the GSM8K dataset (Cobbe et al., 2021) which contains grade-level math word problems that some LLMs struggle with. We also test on the AGIEval-Math dataset (Zhong et al., 2023) which contains much harder questions. The results in Table 4 are based on experiments using GPT-4o-mini and report accuracy as the evaluation metric, rather than F1 score.
The results presented in Table 4 indicate that MAPO’s optimized prompt yields modest accuracy improvements in mathematical reasoning tasks. Specifically, it achieves a 1.2% increase on AGIEval-Math and a 1.0% increase on GSM8K.